Give the exact and approximate solutions to three decimal places.
y^2 - 10y + 25 = 36

Answer by stanbon(75887)

Give the exact and approximate solutions to three decimal places.
y^2 - 10y + 25 = 36
------
y^2-10y-11 = 0
(y-11)(y+1) = 0
-----
y = 11 or y = -1
====================
Cheers,
Stan H.